Doug Orleans

882 Broadway #3R Somerville, MA 02144

dougorleans@gmail.com

http://steak.place.org/dougo/resume/ http://www.linkedin.com/in/dougo http://github.com/dougo

Employment

Jun. 2012–Feb. 2015 PayPal, Boston, MA Tech lead for the team that developed Campaign Manager, a Ruby on Rails web application used by internal customers for managing offers and ads for demand generation.

Dec. 2011–May 2012 MCNA Dental, Fort Lauderdale, FL Developed and maintained various features of DentalTrac, a Ruby on Rails web application used by the entire company for managing dental insurance claims, subscribers, and providers.

Jun. 2009–Dec. 2011 Appleseed Software Consulting, Somerville, MA

As an offsite contract consultant, implemented web applications for clients using the Catalyst framework in Perl, along with MySQL, Template Toolkit, FormFu, and jQuery.

Jul. 2010–Sep. 2010 Art+Culture Editions, New York, NY

As an offsite contract consultant, helped implement an e-commerce web application with Ruby on Rails and Spree.

Jan. 2005–Nov. 2007 Gensym Corporation, Burlington, MA

Member of the G2 Core development team. Maintained and added features to a 20-year-old codebase of over one million lines of Common Lisp. Designed and implemented client and server libraries for HTTP, SOAP, WSDL. Helped implement the G2 Graphical Language, based on WS-BPEL.

1996–2003 Research assistant, Northeastern University, Boston, MA

Member of Dr. Karl Lieberherr’s Demeter research team. Designed, implemented, and maintained DemeterJ, an extension to the Java language for adaptive programming (an early form of aspect-oriented programming). Later, took over development of DJ (Demeter in pure Java) and DAJ (Demeter for AspectJ) from MS student projects.

1992–1995 Software developer, Pure Software, Sunnyvale, CA

Member of the development team for Purify, a memory corruption and leak detection tool for C/C++ programs in Solaris. Worked on the reporting back-end, GUI, and common library code for Purify, PureLink, and Pure-Coverage products.

Education

1995–2005 Ph.D. Computer Science, Northeastern University, Boston, MA

Designed and implemented the Socrates programming language (embedded in PLT Scheme), which unifies and generalizes object-oriented and aspect-oriented programming language mechanisms for separation of concerns using predicate dispatching and open classes.

1988–1992 B.A. Computer Science, University of California, Berkeley, CA

Member of the eXperimental Computing Facility (XCF). Re-implemented the Berkeley Logo interpreter as an explicit-control evaluator rather than a meta-circular evaluator.